OpenStudy (anonymous):

find the 11th term of the arithmetic sequence. a2=-133 and a3=-121

OpenStudy (anonymous):

a + d = -133 a + 2d = -121

OpenStudy (anonymous):

so d = 12 a + 12 = -133 a = -145 t11 = -145 + 10(12)

OpenStudy (anonymous):

term 11 is -25

OpenStudy (anonymous):

Thank you!! Can you help me with my other question?